OpenStudy (anonymous):

Evaluate y' at the point (-1,2) for: (3(x^3)y)-2(x^2)=8

OpenStudy (anonymous):

I keep thinking I have it but keeps getting marked as wrong :/

OpenStudy (anonymous):

Does it simplify to \[y'=\frac{ 4x-9x ^{2} y}{ 3x ^{3}y }\] ?

OpenStudy (anonymous):

is the answer -2/9?

OpenStudy (anonymous):

I don't know the answer. I got 11/3 but it said that was wrong

OpenStudy (anonymous):

i just tried -2/9 also, that didn't work :/

OpenStudy (anonymous):

okay well i put \[3x ^{3}y - 2x ^{2}\] is that right?

OpenStudy (anonymous):

correct

OpenStudy (gorv):

\[3x^3y-2x^2\]

OpenStudy (gorv):

\[9x^2*y+3x^3*dy/dx -4x=0\]

OpenStudy (anonymous):

okay well deriving the first part i got \[3x ^{3}\frac{ dy }{ dx } + 9x ^{2}y\] using product rule

OpenStudy (gorv):

\[3x^2*dy/dx=4x-9x^2*y\]

OpenStudy (gorv):

\[dy/dx =(4x-9x^2*y)/3x^2\]

OpenStudy (anonymous):

\[(3(x^3)y)-2(x^2)=8 \rightarrow (3(x^3)y)=8+2(x^2)\] \[y = \frac {8+2x^2}{3x^3}\] differentiating w.r.t x we find: \[y' = \frac {(3x^3) \frac{d}{dx}(8+2x^2) -(8+2x^2) \frac{d}{dx}(3x^3) }{(3x^3)^2}\] \[y' = \frac {(3x^3)(0+2 \times 2x) -(8+2x^2) \(3 \times 3 x^2) }{(3x^3)^2}\] \[y' = \frac {(3x^3)(4x) -(8+2x^2) (9 x^2) }{(3x^3)^2}\] \[y' = \frac {12x^4 -(72x^2+18x^4)}{(3x^3)^2}\] \[y' = \frac {12x^4 -72x^2-18x^4}{(3x^3)^2}\] \[y' = \frac { -72x^2-6x^4}{9x^6}\] \[y' = \frac { -6x^2(12+x^2)}{9x^6}\] \[y' = \frac { -2(12+x^2)}{3x^4}\] \[y'_{(-1, 2)} = \frac { -2(12+(-1)^2)}{3(-1)^4}= \frac { -2(12+1)}{3}\] \[= \frac { -2(13)}{3}= \frac { -26}{3}\] is the required answer @SRoyea

OpenStudy (anonymous):

then you do the easy \[-4x = 0\] solve for dy/dx \[\frac{ dy }{ dx }=\frac{ 4x-9x ^{2}y }{ 3x ^{3} }\]

OpenStudy (anonymous):

plug in and you get -14/3

OpenStudy (anonymous):

Wow you guys are all really great thanks! Helped so much, got the answer and it makes sense to me now. The answer was 22/3. dpasingh, thanks for taking the time to break it down that far!

OpenStudy (anonymous):

Thanks guys!!